How long does it take an automobile traveling in t
Answer #1

relative speed , v = 50 - 35 = 15 km/h

v = 15 x 1000 m / 3600s = 4.17 m/s

Relative distance have to cover, d= 55 m

t = d/v

= 55/4.17 = 13.2 sec